If consumption is C=100+0.75Yd

Taxes is T=50+0.5Y

Export is X=200

Import is M=50+0.25Y

Government spending is G=150

Investment is I=200. Y is domestic income and Y is private disposable income. Determine the trade balance at equilibrium


1
Expert's answer
2021-05-16T17:49:41-0400

Consumption is C=100+0.75Yd

Taxes is T=50+0.5Y

Export is X=200

Import is M=50+0.25Y

Government spending is G=150

Investment is I=200

here;

Y is domestic income

Yd= Private disposable income


"Y= C+I+G+(X\u2212M)"

"Y= 100+0.75(Y-T)+200+150+(200\u2212(50+0.25Y))"

"Y=100+0.75(Y\u2212(50+0.5Y)+350+(200-50-0.25Y)"

"Y=100+0.75(Y-50-0.5Y)+350+150-0.25Y"

"Y=100+0.75(0.5Y\u221250)+500\u22120.25Y"

"Y=600+0.375Y\u221237.5\u22120.25Y"

"Y=562.5+0.125Y"

"Y\u22120.125Y=562.5"

"0.875Y=562.5"

"Y=642.86"


"M=50+0.25Y"

"M=50+0.25(642.86)"

"M=50+160.715"

"M=210.715"

"Trade \space Balance= Export−Import"

           "= 200\u2212210.715"

           "= \u221210.715"

Since the trade balance is negative 10.714 that is,. there is trade deficit in the economy.